# Large deviations in non-Markovian stochastic epidemics

### v1

@@ Abstract
The QSD can be explicitly written using Eq. (5) in terms of the $_{2}F_{1}(a,b;c;z)$ hypergeometric function, and generalizes the result for Markovian reactions [20, 23]. In Fig. 3 we show excellent agreement between simulated QSDs for various values of $\alpha$ and our theoretical prediction. We also plot by dashed lines the QSDs that would be obtained if we use a Markovian theory with adjusted reaction rates, obtained by replacing the gamma-distributed recovery process with an exponential one, and adjusting the infection rate by $R_{0}\to R_{0}/R_{bif}$. While being a good approximation close to the mean, the tails are clearly missed by this adjusted distribution, which demonstrates that the non-Markovian effects cannot be fully captured by simply adjusting the mean of a Markovian reaction.
We can also compute the QSD’s standard deviation $\sigma$ by approximating $\pi(x_{i})$ as a Gaussian around $x_{i}^{*}$ [23], which yields $\sigma=\left|-N\mathcal{S}^{\prime\prime}(x_{i}^{*})\right|^{-1/2}$. Using Eq. (5) we find
$\sigma\!=\!\sqrt{2^{1/\alpha\!-\!1}/(NR_{0})}\!=\!\sqrt{(1\!-\!x_{i}^{*})/[2\alpha N(1\!-\!2^{-1/\alpha})}.$ (13)
First of all, at $\alpha=1$ we recover the Markovian result of $\sigma=(NR_{0})^{-1/2}$. Second, as $\alpha$ decreases below $1$, the standard deviation, or typical fluctuations, can greatly increase, making disease extinction much more likely, see below. In fact, we can approximate $\alpha$ close to $1$ to obtain $\sigma=(NR_{0})^{-1/2}\left[1+\ln{2}/2(1-\alpha)\right]$, from which it is clear that as $\alpha$ decreases below $1$, the variance grows.
Figure 3: QSDs at $\alpha\!=\!0.45,0.6,1,10$ in (a)–(d), for Markovian infection and gamma-distributed recovery, with N=1000N\!=\!1000, $R_{0}\!=\!2$ and $10^{5}$ runs per $\alpha$. Here we compare simulations (symbols) to theory (solid line) and to a Markovian theory adjusted to the same mean (dashed line), with $x_{i}(0)=x_{i}^{*}$.
@@
In Fig. 4(a,b) we show how the shape parameter $\alpha$ affects the metastable mean (11) and its standard deviation (13), both normalized by their Markovian values ($x_{i}^{*}(1)=1/3$, $\sigma(1)\approx 0.0258$ for $R_{0}=1.5$). As $\alpha$ decreases, the mean drastically decreases and the variance increases; this is since the typical time between recovery events gets shorter as the WT distribution becomes wider and more skewed. The opposite occurs for $\alpha>1$; for $\alpha\to\infty$, the mean approaches $x_{i}^{*}(\infty)/x_{i}^{*}(1)\simeq(1-\ln 2/R_{0})/(1/3)\simeq 1.614$. Near the bifurcation, the mean vanishes and the WKB approximation breaks down.
@@
The dependence of the variance on $\alpha$ as observed in Fig. 4(b) can be explained by looking at the right-hand-side of Eq. (13), where the standard deviation is expressed as a function of the mean $x_{i}^{*}$ and $\alpha$, assuming the effect of non-Markovianity on the mean is corrected by adjusting $R_{0}$  [^2] Notably, the dependence of $\sigma$ on $x_{i}^{*}$ and $\alpha$ remains identical also when infection is gamma-distributed and recovery is exponential, even though $x_{i}^{*}$ changes in that case.. This reveals three effects shaping $\sigma$. The first and strongest is the monotone relation $\sigma\sim\sqrt{1-x_{i}^{*}}$, dominating the behavior. The second stems from non-Markovianity, $\sigma\sim\left[2\alpha(1-2^{-1/\alpha})\right]^{-1/2}$; it reinforces the main effect in the non-Markovian recovery case of increase of $\sigma$ as $\alpha$ is decreased (but partially opposes it for non-Markovian infection, see SI, Sec. C). The third arises from WKB breakdown: as $x_{i}^{*}$ approaches 0, the monotone relation fails, and $\sigma$ begins to decrease with $x_{i}^{*}$, as fluctuations near 0 are constrained by the boundary (similarly as when $x_{i}^{*}$ approaches $1$).
Finally, having computed the complete QSD (12), we can evaluate the MTE, using $\tau_{ext}\sim e^{N\mathcal{S}(0)}$ [54, 23, 25], which gives rise to an explicit expression containing the hypergeometric function $_{2}F_{1}(a,b;c;z)$. A simplified expression can be obtained, in the limit of $|\alpha-1|\ll 1$, which can provide insight on how the disease lifetime changes as one slightly deviates from Markovianity. Approximating $\mathcal{S}(0)$ in the leading order in $|\alpha-1|\ll 1$, we find
$\mathcal{S}(0)$ $\simeq$ $\ln{R_{0}}+R_{0}^{-1}-1-\left\{\ln(R_{0}\!+\!1)-2\right$.
$\hskip-31.29802pt\!-\!\left.R_{0}^{-1}\!\left[\pi^{2}/12\!-\!2\!+\!2\ln{2}\!-\!\ln(R_{0}\!+\!1)\!+\!\text{Li}_{2}(-R_{0})\right]\right\}\!(1-\alpha),$
where $\text{Li}_{2}(z)$ is the second polylogarithm. Note that we recover the known result of $\mathcal{S}(0)=1-\ln{R_{0}}-1/R_{0}$ at $\alpha=1$ [20, 22, 23, 25]. In Fig. 4(c) we plot the MTE versus $\alpha$ for exponential infection and gamma-distributed recovery. Good agreement holds as long as the action is large (i.e., $\ln\tau_{ext}\gg 1$). We also show the approximate theoretical result close to $\alpha=1$, which excellently agrees with simulations well beyond its region of applicability.
@@
Figure 4: Normalized mean $x_{i}^{*}$ (a) and normalized standard deviation $\sigma$ (b) vs $\alpha$, for Markovian infection and gamma-distributed recovery, $N=1000$, $R_{0}=1.5$, and $10^{3}$ runs per $\alpha$. Inset zooms in close to the bifurcation. Simulations (symbols) are compared with theory (solid line). In (c) we plot the MTE vs $\alpha$ for the same WTs, $N=180$, $R_{0}=1.5$, and $10^{2}$ runs per $\alpha$. Simulations (symbols) are compared with theory (solid line) and Eq. (Large deviations in non-Markovian stochastic epidemics) (dashed line). In all panels $x_{i}(0)=x_{i}^{*}$.
Conclusion. We studied how non-Markovian infection and recovery WTs shape long-term epidemic dynamics. For both the SIR and SIS models, we derived an effective asymptotic master equation, valid at late times, by explicitly computing the memory kernels given the WT distributions. Within the SIR model we showed that non-Markovian recovery can drastically increase the final outbreak size distribution’s width, which may have key implications on pandemics handling. Within the SIS model, we computed the QSD of the long-lived metastable state and the MTE, and showed that the latter can greatly decrease as memory strength increases, making epidemics far more fragile to extinction. Finally, we checked that other WT distributions (e.g., power laws) give rise to qualitatively similar behavior, see SI, Sec. D.
In future work, it is desirable to compare our predictions to realistic non-Markovian epidemic scenarios with empirical WT distributions. In addition, while mean-field impacts of non-Markovian epidemics on structured populations were recently analyzed, stochastic fluctuations remain largely unexplored. Capturing them will likely require a modified formalism to handle non-resetting reactions that are able to account for memory effects in heterogeneous contact structures.
@@
We begin by showing how the non-Markovian infection and recovery processes can be formulated. In this aspect, the analysis of the SIS and SIR models will be identical, but later the treatment of the master equation will be different depending on the model at hand. We start by computing the survival probabilities corresponding to the WT distributions defined in the main text
$\Psi_{1}(t)=e^{-\lambda_{1}t},\quad\Psi_{2}(t)=\frac{\Gamma(\alpha,\alpha\lambda_{2}t)}{\Gamma(\alpha)}.$ (S1)
We next calculate $\phi_{1}(t)$—the distribution for a single infection event to occur at time $t$, assuming no recovery has occurred up to $t$, and $\phi_{2}(t)$—the distribution for a single recovery event to occur at time $t$, assuming no infection has occurred up to $t$. This yields $\phi_{1}(t)=\psi_{1}(t)\Psi_{2}(t)$, $\phi_{2}(t)=\psi_{2}(t)\Psi_{1}(t)$, and with Eq. (2) in the main text we get
$\phi_{1}(t)=NR_{0}x_{s}x_{i}\frac{\Gamma(\alpha,\alpha Nx_{i}t)}{\Gamma(a)}e^{-NR_{0}x_{s}x_{i}t},$
$\phi_{2}(t)=\frac{(\alpha Nx_{i})^{a}t^{a-1}e^{-\alpha Nx_{i}t}}{\Gamma(a)}e^{-NR_{0}x_{s}x_{i}t}.$ (S2)
where $\Gamma\left(a,z\right)=\int_{z}^{\infty}t^{z-1}e^{-t}dt\,$ is the incomplete gamma function. We then calculate the Laplace transform of $\tilde{\phi}_{i}$:
@@
Finally, we use the finite value theorem $\lim_{t\to\infty}f(t)=\lim_{u\to 0}u\tilde{f}(u)$, and define the normalized asymptotic memory kernel $\mathcal{M}_{i}$ as: $\mathcal{M}_{i}:=(1/N)\lim_{t\to\infty}M_{i}(t)=(1/N)\lim_{u\to 0}\tilde{M}_{i}(u)$ to get Eq. (4) and Eq. (5) in the main text.
II B. Derivation of the final outbreak-size distribution
@@
$\hskip-25.60747pt\dot{x}_{s}=\frac{\partial\mathcal{H}}{\partial p_{s}}=-\mathcal{M}_{1}e^{p_{i}-p_{s}},$ (S7)
$\hskip-25.60747pt\dot{x}_{i}=\frac{\partial\mathcal{H}}{\partial p_{i}}=\mathcal{M}_{1}e^{p_{i}-p_{s}}-\mathcal{M}_{2}e^{-p_{i}},$ (S8)
$\hskip-25.60747pt\dot{p}_{s}=\!\!-\frac{\partial\mathcal{H}}{\partial x_{s}}\!=\!\frac{\partial\mathcal{M}_{1}}{\partial x_{s}}\left(1\!-\!e^{p_{i}-p_{s}}\right)\!+\!\frac{\partial\mathcal{M}_{2}}{\partial x_{s}}\left(1\!-\!e^{-p_{i}}\right)\!\!,$ (S9)
$\hskip-25.60747pt\dot{p}_{i}=\!\!-\frac{\partial\mathcal{H}}{\partial x_{i}}\!=\!\frac{\partial\mathcal{M}_{1}}{\partial x_{i}}\left(1\!-\!e^{p_{i}-p_{s}}\right)\!+\!\frac{\partial\mathcal{M}_{2}}{\partial x_{i}}\left(1\!-\!e^{-p_{i}}\right)\!\!.$ (S10)
@@
Before delving into these equations we point out an important constant of motion, which appears in the SIR model and enables the intrinsically two-dimensional problem to be cast into one dimension. First, we point out that the Hamiltonian is a constant of motion, since the rates do not depend on time explicitly. Furthermore, when examining the rates $\mathcal{M}_{i}$ we can see that in our example case they are both linear in $x_{i}$, see Eq. (5) in the main text. This attribute also holds for a wide variety of other WT distributions such as a power-law distribution, see SI Sec. D. As a result, using Eqs. (S6) and (S10), we can write the Hamiltonian in a suggestive way $\mathcal{H}=-x_{i}\dot{p_{i}}$. Finally, since most epidemic waves start from a small fraction of infected, we argue that typically $x_{i}(t=0)\ll 1$, and is certainly not macroscopic; in fact, it is often assumed that one starts with one infected individual such that $x_{i}(t=0)\sim O(1/N)$ with $N\gg 1$. As a result, since $\mathcal{H}(t=0)$ is a constant of motion, one must have $\mathcal{H}\simeq 0$ throughout the epidemic. Yet, due to the fact that $\mathcal{H}=-x_{i}\dot{p_{i}}$, and that $x_{i}(t)$ changes from a very small value to a macroscopic fraction of the population during the epidemic wave, the only way the Hamiltonian can stay zero is by having $\dot{p_{i}}=0$ throughout it. Therefore, we have found a new constant of motion: $p_{i}$.
Notably, even though in this case $\mathcal{H}\simeq 0$, there is an important distinction between our use of the WKB assumption here and in the SIS model. While the condition of $\mathcal{H}\simeq 0$ is inherent in the SIS model as the distribution becomes metastable with $\partial P(x_{i},t)/\partial t\simeq 0$ (up to exponential accuracy), in the SIR model this condition solely stems from starting with a small fraction of infected individuals.
We now derive the outbreak size distribution by adopting the methodology of [27] and defining a new constant of motion $m\!=\!e^{p_{i}}$. Using Eqs. (S7) and  (S8) and the fact that the total population is constant, i.e. $\dot{x}_{s}+\dot{x}_{i}+\dot{x}_{r}=0$, we get an equation for $\dot{x}_{r}$
$\dot{x}_{r}=-\dot{x}_{s}-\dot{x}_{i}=\mathcal{M}_{2}/m.$ (S11)
By equating the Hamiltonian [Eq. (S6)] to zero, we can get an expression for $e^{p_{s}}$
@@
Now we divide Eq. (S7) by (S11) to obtain a differential equation for the dependence of $x_{s}$ on $x_{r}$
@@
By integrating $x_{s}$ from $1$ to $x_{s}^{*}$ and $x_{r}$ from 0 to $1-x_{s}^{*}$, we obtain an implicit relation between the final outbreak size $x_{s}^{*}=x_{s}(t\to\infty)$ and $m$
@@
Note that, plugging $\alpha=1$ into $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ and using Eq. (5) in the main text, this result coincides with that of Ref. [27] in the Markovian case. Equation (S14) is important since it removes $m$ from the action $\mathcal{S}$ and makes it a function of a single variable $x_{s}^{*}$. To find the action $\mathcal{S}(x_{s}^{*})$ for each final state $x_{s}^{*}$, which determines the outbreak size distribution via the WKB ansatz, we write down the formal solution of the action, as done in classical mechanics [23,25,27]

### v2

@@ Abstract
The QSD can be explicitly written using Eq. (5) in terms of the elementary functions, for each value of $\alpha$, and generalizes the result for Markovian reactions [48, 7]. In Fig. S1 we show excellent agreement between simulated QSDs for various values of $\alpha$ and our theoretical prediction, see details in the SI, Sec. D.
We can also compute the QSD’s variance, $\sigma^{2}$, via a Gaussian approximation around $x_{i}^{*}$ [7]. This yields
${\color[rgb]{0,0,0}\sigma^{2}=\frac{2^{1/\alpha-1}}{NR_{0}\alpha^{2}(2^{1/\alpha}-1)^{2}}}=\frac{1-x_{i}^{*}}{2\alpha N(1-2^{-1/\alpha})},$ (14)
@@
where we have used Eq. (5). At $\alpha=1$, the Markovian result is recovered, $\sigma^{2}\!=\!1/(NR_{0})$. As $\alpha$ increases, typical fluctuations can greatly increase, making disease clearance more likely. In Fig. 3(a,b) we show how the shape parameter $\alpha$ affects the metastable mean (12) and its standard deviation (14), both normalized by their Markovian values $x_{i}^{*}(\alpha\!=\!1)\!=\!1/3$, $\sigma(\alpha\!=\!1)\!\simeq\!0.026$ for $R_{0}=1.5$]. As $\alpha$ increases, the mean decreases drastically [60, 13], and the variance increases; this occurs since the typical time between infection events gets longer as the WT distribution becomes narrower. At $\alpha\to 0$, the times between infection events is very short and the mean approaches 11, while the variance approaches 0.
@@
The dependence of the variance on $\alpha$ as observed in Fig. 3(b) can be explained by looking at the right-hand-side of Eq. (14), where the standard deviation is expressed as a function of the mean $x_{i}^{*}$ and $\alpha$ (instead of $R_{0}$ and $\alpha$ [^5] Notably, the dependence of $\sigma$ on $x_{i}^{*}$ and $\alpha$ remains identical also when infection is exponential and recovery is gamma-distributed, even though $x_{i}^{*}$ changes in that case.. This reveals two effects shaping $\sigma$. The first and strongest is the monotone relation $\sigma\!\sim\!\sqrt{1\!-\!x_{i}^{*}}$, dominating the behavior. The second effect stems from non-Markovianity, $\sigma\!\sim\!\left[2\alpha(1\!-\!2^{-1/\alpha})\right]^{-1/2}$; for non-Markovian infection, it partially opposes the main effect of increase in $\sigma$ as $\alpha$ is increased; yet, it reinforces the main effect in the case of non-Markovian recovery (see SI, Sec. E). Notably, as $x_{i}^{*}$ approaches 0, the WKB approximation is expected to break down as the action is no longer large. This effect is more evident for non-Markovian recovery, see SI, Sec. E.
Finally, having computed the complete QSD (13), we can evaluate the MTE, using $\tau_{ext}\sim e^{N\mathcal{S}(0)}$, where $\mathcal{S}(0)$ is the action barrier to extinction [20, 7, 9, 8]. The integral in Eq. (13) from $x_{i}^{*}$ to 0 yields the MTE, and can be explicitly computed for each $\alpha$. In particular, a simplified expression can be obtained, at $|\alpha-1|\ll 1$, which can provide insight on how the disease lifetime changes as one deviates from Markovianity. In the leading $|\alpha-1|\ll 1$ order, we obtain $\mathcal{S}(0)\!\simeq\!{\cal S}_{0}\!-\!(\alpha\!-\!1)/(2R_{0})\!\!\left[\!(R_{0}\!+\!1)^{2}\ln\!\left(\!1\!+\!R_{0}^{-1}\!\right)\!-\!R_{0}\!+\!1\!-\!\ln(R_{0}/16)\right]$, where ${\cal S}_{0}=\ln R_{0}+R_{0}^{-1}-1$ is the action barrier in the $\alpha=1$ case [48, 47, 7, 8]. In Fig. 3(c) we plot the MTE versus $\alpha$ for gamma-distribued infection and exponential recovery. Very good agreement holds as long as the action is large (i.e., $\ln\tau_{ext}\gg 1$), which is the case for the entire range of the figure. Here, the solid line shows the theoretical result, $\tau_{ext}\simeq A(\alpha,R_{0},N)e^{N\mathcal{S}(0)}$; it includes an additional preexponential factor $A(\alpha,R_{0},N)$ [7, 8], numerically fitted (for fixed NN and $R_{0}$) to be $A\sim\alpha^{2}$.
@@
Figure 3: In (a) and (b) we show the normalized mean $x_{i}^{*}$ and normalized standard deviation $\sigma$ versus $\alpha$, for gamma-distributed infection and Markovian recovery. Here, $N=1000$, $R_{0}=1.5$, and we ran $10^{3}$ realizations per $\alpha$. Simulations (symbols) are compared with theory (line). In (c) we plot the MTE versus $\alpha$ for the same WTs, $N=110$, $R_{0}=1.51$ and $10^{2}$ runs per $\alpha$. Simulations (symbols) are compared with theory (line), see text. In (a-c), $x_{i}(0)=x_{i}^{*}$.
Discussion. We have studied how non-Markovian reactions shape long-term epidemic dynamics, by deriving the late-time asymptotics of the master equation, and memory kernels emanating from the WTs. We have shown, within the SIR model, that under non-Markovian infection and exponential recovery, the mean outbreak size and its entire distribution strongly depend on the shape parameter $\alpha$; notably, as $\alpha$ grows, the outbreak risk greatly diminishes. Within the SIS model, we have shown that as $\alpha$ increases, disease prevalence decreases [60, 13], and the risk of disease eradication greatly increases.
Having focused so far on non-Markovian infection, we now show its applicability to realistic scenarios with both infection and recovery being non-Markovian. Empirical studies of acute infections like influenza and COVID-19 consistently report nonexponential (Gamma, Weibull, or log-normal) WTs for infection and recovery, typically with gamma shape parameters $1<\alpha_{\text{inf}},\alpha_{\text{rec}}<6$ [69, 12, 42, 24, 1, 49, 40, 16, 14, 64]. Infection WTs typically measure the time from inferred exposure or symptom onset to secondary transmission, while recovery WTs correspond to infectious periods or time from a positive test to a negative one. Population-level heterogeneity further shapes these distributions. For example, it is expected that elderly, immunocompromised, or high-contact individuals will show longer, more variable recovery and lower shape parameters, whereas children and young adults in community settings will tend to progress and recover more uniformly [1, 40, 64].
In Fig. 4 we show the dramatic impact of incorporating non-Markovian reactions. Here numerical and theoretical heatmaps of the outbreak size distribution’s coefficient of variation (COV) are compared for gamma-distributed infection and recovery, versus the shape parameters in a well-mixed setting. Notably, the COV, indicating the relative error in the expected outbreak size, can greatly exceed the Markovian prediction, especially in regions of large infection shape parameter, indicating synchronized and rapid progression from exposure to infectiousness.
Figure 4: Numerical (left) and theoretical (right) heatmaps of the outbreak-size distribution COV (normalized by the Markovian COV) versus the infection and recovery shape parameters, for a well-mixed system with N=5000N\!=\!5000, $R_{0}\!=\!2$, and $10^{3}$ runs per point. Black lines are equi-COV contour lines.
Future research should aim at extending our formalism to realistic non-Markovian epidemic settings involving structured populations residing on complex heterogeneous networks. The mean-field aspects of non-Markovian epidemics have already been studied on such networks, and the next challenge is to study the interplay between large deviations and complex network topology under non-Markovian dynamics. Therefore, upon successful extension of the theory to real-life network topologies, and accounting for empirically grounded shape parameters for both infection and recovery (as done in Fig. 4), this framework is expected to provide a more accurate representation of epidemic dynamics, improving estimates of epidemic thresholds, outbreak sizes, disease clearance times, and most importantly, intervention measures.
Acknowledgments. The authors wish to thank Ami Taitelbaum for many useful discussions.
@@
We begin by showing how the non-Markovian infection and recovery processes can be formulated. In this aspect, the analysis of the SIS and SIR models are identical, but later the treatment of the master equation will be different depending on the model at hand. We start by computing the survival probabilities corresponding to the waiting-time (WT) distributions between infection and recovery events. For simplicity, we outline the derivation for the case of Markovian infection and non-Markovian recovery for which the WTs are given by Eq. (S21) below. Yet, the derivation for the complementary case is identical. The corresponding survival probabilities are
$\Psi_{1}(t)=e^{-\lambda_{1}t},\quad\Psi_{2}(t)=\frac{\Gamma(\alpha,\alpha\lambda_{2}t)}{\Gamma(\alpha)},$ (S1)
where $\Gamma\left(a,z\right)=\int_{z}^{\infty}t^{a-1}e^{-t}dt\,$ is the upper incomplete gamma function. We next calculate $\phi_{1}(t)$—the distribution for a single infection event to occur at time $t$, assuming no recovery has occurred up to $t$, and $\phi_{2}(t)$—the distribution for a single recovery event to occur at time $t$, assuming no infection has occurred up to $t$. This yields $\phi_{1}(t)=\psi_{1}(t)\Psi_{2}(t)$, $\phi_{2}(t)=\psi_{2}(t)\Psi_{1}(t)$, and with Eq. (S21) below we get
$\phi_{1}(t)$ $=$ $NR_{0}x_{s}x_{i}\frac{\Gamma(\alpha,\alpha Nx_{i}t)}{\Gamma(\alpha)}e^{-NR_{0}x_{s}x_{i}t},$
$\phi_{2}(t)$ $=$ $\frac{(\alpha Nx_{i})^{\alpha}t^{\alpha-1}e^{-\alpha Nx_{i}t}}{\Gamma(\alpha)}e^{-NR_{0}x_{s}x_{i}t}.$ (S2)
We then calculate the Laplace transform of $\tilde{\phi}_{i}$:
@@
Finally, we use the finite value theorem $\lim_{t\to\infty}f(t)=\lim_{u\to 0}u\tilde{f}(u)$, and define the normalized asymptotic memory kernel $\mathcal{M}_{i}$ as: $\mathcal{M}_{i}:=(1/N)\lim_{t\to\infty}M_{i}(t)=(1/N)\lim_{u\to 0}\tilde{M}_{i}(u)$ to get Eq. (4) with effective memory kernels: $\mathcal{M}_{1}=R_{0}x_{i}x_{s}$ and $\mathcal{M}_{2}=R_{0}x_{i}x_{s}/\left[(1+R_{0}x_{s}/\alpha)^{\alpha}-1\right]$. These results complement Eq. (5) in the main text.
II B. Different choices of waiting-time distributions
In this section, we show that the asymptotic memory kernels can be obtained in a similar manner also for other choices of WT distributions; for concreteness, we focus on the case of a power-law inter-event time distribution. To simplify matters, we consider here the complementary choice of Markovian infection and non-Markovian recovery, where in Fig. 4 in the main text and in Fig. S3 we show results when both infection and recovery are non-Markovian.
Under exponential infection and power-law distributed recovery, WT distributions are:
with means of $\lambda_{1}=NR_{0}x_{s}x_{i}$ and $\lambda_{2}=Nx_{i}$. Here, the shape parameter is bounded, $\alpha>1$, to ensure a finite mean, and in the limit $\alpha\to\infty$ the distribution converges to an exponential one. Notably, the regime of $\alpha>1$ in the gamma distribution does not exist here; namely, by choosing a power-law WT we can only increase the distribution’s width compared to an exponential one. Performing the explicit calculation of the asymptotic memory kernels, we find
where $E_{m}(z)\equiv\int_{1}^{\infty}e^{-z\tau}\tau^{-m}d\tau\,$ is the exponential integral function.
At this point, one can plug these memory kernels [Eqs. (S22) and (S7)] into the late-time master equations and perform calculations along the same lines as in the main text, which allows to find the quantities of interest, e.g., outbreak-size distribution within the SIR model, or MTE within the SIS model.
III C. Derivation of the final outbreak-size distribution
@@
$\hskip-19.91692pt\dot{x}_{s}$ $=$ $\frac{\partial\mathcal{H}}{\partial p_{s}}=-\mathcal{M}_{1}e^{p_{i}-p_{s}},$ (S9)
$\hskip-19.91692pt\dot{x}_{i}$ $=$ $\frac{\partial\mathcal{H}}{\partial p_{i}}=\mathcal{M}_{1}e^{p_{i}-p_{s}}-\mathcal{M}_{2}e^{-p_{i}},$ (S10)
$\hskip-19.91692pt\dot{p}_{s}$ $=$ $-\frac{\partial\mathcal{H}}{\partial x_{s}}\!=\!\frac{\partial\mathcal{M}_{1}}{\partial x_{s}}\left(1\!-\!e^{p_{i}-p_{s}}\right)\!+\!\frac{\partial\mathcal{M}_{2}}{\partial x_{s}}\left(1\!-\!e^{-p_{i}}\right)\!,$ (S11)
$\hskip-19.91692pt\dot{p}_{i}$ $=$ $-\frac{\partial\mathcal{H}}{\partial x_{i}}\!=\!\frac{\partial\mathcal{M}_{1}}{\partial x_{i}}\left(1\!-\!e^{p_{i}-p_{s}}\right)\!+\!\frac{\partial\mathcal{M}_{2}}{\partial x_{i}}\left(1\!-\!e^{-p_{i}}\right)\!.$ (S12)
@@
Before delving into these equations we point out an important constant of motion, which appears in the SIR model and enables the intrinsically two-dimensional problem to be cast into one dimension. First, we point out that the Hamiltonian is a constant of motion, since the rates do not depend on time explicitly. Furthermore, when examining the rates $\mathcal{M}_{i}$ we can see that in our example case they are both linear in $x_{i}$, see Eq. (5) in the main text. This attribute also holds for a wide variety of other WT distributions such as a power-law distribution, see SI Sec. B. As a result, using Eqs. (S8) and (S12), we can write the Hamiltonian in a suggestive way $\mathcal{H}=-x_{i}\dot{p_{i}}$. Finally, since most epidemic waves start from a small fraction of infected, we argue that typically $x_{i}(t=0)\ll 1$, and is certainly not macroscopic; in fact, it is often assumed that one starts with one infected individual such that $x_{i}(t=0)\sim O(1/N)$ with $N\gg 1$. As a result, since $\mathcal{H}(t=0)$ is a constant of motion, one must have $\mathcal{H}\simeq 0$ throughout the epidemic. Yet, due to the fact that $\mathcal{H}=-x_{i}\dot{p_{i}}$, and that $x_{i}(t)$ changes from a very small value to a macroscopic fraction of the population during the epidemic wave, the only way the Hamiltonian can stay zero is by having $\dot{p_{i}}=0$ throughout it. Therefore, we have found a new constant of motion: $p_{i}$.
Notably, even though in this case $\mathcal{H}\simeq 0$, there is an important distinction between our use of the WKB assumption here and in the SIS model. While the condition of $\mathcal{H}\simeq 0$ is inherent in the SIS model as the distribution becomes metastable with $\partial P(x_{i},t)/\partial t\simeq 0$ (up to exponential accuracy), in the SIR model this condition solely stems from starting with a small fraction of infected.
We now derive the outbreak size distribution by adopting the methodology of [28] and defining a new constant of motion $m\!=\!e^{p_{i}}$. Using Eqs. (S9) and  (S10) and the fact that the total population is constant, i.e. $\dot{x}_{s}+\dot{x}_{i}+\dot{x}_{r}=0$, we obtain $\dot{x}_{r}=-\dot{x}_{s}-\dot{x}_{i}=\mathcal{M}_{2}/m$. By equating the Hamiltonian [Eq. (S8)] to zero, we can get an expression for $e^{p_{s}}$
@@
Now we divide $\dot{x}_{s}$ from Eq. (S9) by $\dot{x}_{r}=\mathcal{M}_{2}/m$ to obtain a differential equation for the dependence of $x_{s}$ on $x_{r}$
@@
By integrating $x_{s}$ from 11 to $x_{s}^{*}$ and $x_{r}$ from 0 to $1-x_{s}^{*}$, we obtain an implicit relation between the final outbreak size $x_{s}^{*}=x_{s}(t\to\infty)$ and $m$
@@
Note that, plugging $\alpha=1$ into $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ and using Eq. (5) in the main text, this result coincides with that of Ref. [28] in the Markovian case. Equation (S15) is important since it removes $m$ from the action $\mathcal{S}$ and makes it a function of a single variable $x_{s}^{*}$. To find the action $\mathcal{S}(x_{s}^{*})$ for each final state $x_{s}^{*}$, which determines the outbreak size distribution via the WKB ansatz, we write down the formal solution of the action [25,28]